OpenStudy (lilshane):

Draw a number line from –2 to 2. Put tick marks every one-tenth unit. Graph a point with the given coordinate. point N, coordinate –1.7 Which is true about point N on the number line? A. Point N is 0.7 unit to the right of –1. B. Point N is 0.7 unit to the left of –1. C. Point N is 0.7 unit to the left of 0. D. Point N is 0.7 unit to the left of 1.

OpenStudy (lilshane):

Im gona draw the numberline now

OpenStudy (anonymous):

Okay great! Let me know when you're done. You understand how to do the ticks every 1/10 of the way right?

OpenStudy (lilshane):

I just noticed to had to do decimals -_-

OpenStudy (anonymous):

Do you understand how to do that?

OpenStudy (lilshane):

Im gonna try but can you draw one while im drawing one and we can check work together?

OpenStudy (anonymous):

Here i'll explain first and i think you'll get hwo to do it.

OpenStudy (lilshane):

oki

OpenStudy (anonymous):

so you know that 10/10 is a whole number right? Lets think of it as our number 2, even though mathematically 10/10 = 1. so if we start from 0, the first number will be 1/10, 2/10, 3/10.... etc and so on. BUT!!! since 10/10 is equal to 1, we multiply it by 2. 20/10 = 2. This is easy to know because you just cross the zeros out. 2/1 = 2. :D

OpenStudy (anonymous):

so from 0 it goes, 1/10, 2/10,3/10,4/10,5/10,6/10,7/10,8/10,9/10,10/10==1, 11/10,12/10,13/10,14/10,15/10,16/10,17/10,18/10,19/10,20/10==2

OpenStudy (anonymous):

Now how in the world do we know what 1/10th is or 15/10 is? Well in the case of tenth's it is super simple. 1/10 = .1 15/10 = 1.5 We do this by pretending that there is a decimal in the fractions. Let me draw this to show you. forwarning, my hand writign blows.

OpenStudy (anonymous):